### v2

# SubTrack++ : Gradient Subspace Tracking for Scalable LLM Training

Abstract
Training large language models (LLMs) is highly resource-intensive due to their massive number of parameters and the overhead of optimizer states. While recent work has aimed to reduce memory consumption, such efforts often entail trade-offs among memory efficiency, training time, and model performance. Yet, true democratization of LLMs requires simultaneous progress across all three dimensions. To this end, we propose SubTrack++ that leverages Grassmannian gradient subspace tracking combined with projection-aware optimizers, enabling Adam’s internal statistics to adapt to changes in the optimization subspace. Additionally, employing recovery scaling, a technique that restores information lost through low-rank projections, further enhances model performance. Our method demonstrates SOTA convergence by exploiting Grassmannian geometry and achieves lowest evaluation loss, outperforming the current SOTA while reducing pre-training wall time by 43% and maintaining the memory footprint on a 1B-parameter Llama model.
1 Introduction
Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity. However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact and carbon footprint. (Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024).
Several techniques have been proposed to mitigate memory bottlenecks, such as gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020). LoRA (Hu et al., 2021) and other low-rank adaptation methods (Dettmers et al., 2024; Hu et al., 2021; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024) have gained popularity by optimizing a reduced set of parameters. Such approaches often assume that the full parameter space is inherently low-rank—an assumption that does not always hold and can lead to suboptimal performance. In addition, methods like BAdam (Luo et al., 2024) and Block-LLM (Ramesh et al., 2024), utilize block coordinate descent to optimize parameter subsets, achieving memory savings at the cost of partial parameter tuning and reduced accuracy.
Figure 1: We compare baselines on pre-training a 1B-parameter model. (a) SubTrack++ achieves the lowest evaluation loss across all methods. (b) Its peak memory usage is significantly lower than APOLLO and LDAdam, and on par with GaLore and Fira. (c) In terms of wall-time, SubTrack++ incurs minimal overhead relative to APOLLO and is markedly faster than GaLore, Fira, and LDAdam. Overall, SubTrack++ outperforms all baselines in evaluation loss while matching or exceeding them in memory and runtime efficiency.
However, memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024). To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024; Rajabi and Rambhatla, 2024a). Leveraging the fact that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023) , GaLore (Zhao et al., 2024) reduces memory usage by projecting gradients into a low-rank subspace and periodically updating the rank-$r$ approximation via singular value decomposition (SVD)—an approach also used in more recent method, Fira (Chen et al., 2025). While SVD offers optimal low-rank approximation (Robert et al., 2025), performing it repeatedly can pose several challenges. First, it is computationally intensive, and alternatives like random projections (as in APOLLO (Zhu et al., 2025) and GoLore (He et al., 2025)) or approximation methods for estimating dominant singular values (as in LDAdam (Robert et al., 2025), Online Subspace Descent (Liang et al., 2024) and Rajabi and Rambhatla (2024b)) match or outperform SVD in practice. Moreover, SVD is sensitive to noise (Vaswani et al., 2018; He et al., 2025) and tends to degrade in late training stages when gradients are small, often hindering convergence (He et al., 2025).
Geometry-based methods have demonstrated strong performance in various machine learning applications (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Rajabi and Rambhatla, 2024c; Blocker et al., 2023). In this context, using Grassmannian manifolds for subspace tracking has led to structurally embedded information, lower computational complexity, and improved performance (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023; Chakraborty et al., 2017; Zhang and Balzano, 2016). Grassmannian is the set of $r$-dimensional subspaces in a d-dimensional space, and the ultimate goal of low-rank optimization is to find the proper subspace in this space. This line of work has demonstrated robustness and efficiency in high-dimensional, noisy environments (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Balzano et al., 2018; Chakraborty et al., 2017), outperforming online or robust Principle Component Analysis (PCA) techniques (Li, 2004; Balzano et al., 2018; Chakraborty et al., 2017). Their natural robustness against perturbations and strong theoretical guarantees make them particularly well-suited to tracking the evolving gradient subspaces encountered in LLM training. (Balzano, 2012; Balzano et al., 2011).
Figure 2: Visualization of Grassmannian-based subspace tracking: Between subspace updates, gradients are projected onto a fixed subspace. The tangent vector $\nabla F$ is computed via the derivative of a loss function that measures the subspace estimation error. The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$, on the Grassmannian manifold to minimize estimation error.
To this end, we employ subspace tracking on Grassmannian geodesics (Figure 2), to develop a geometry-based, time- and memory-efficient method for tracking the gradient subspace. This way, instead of reconstructing low-rank approximations via expensive SVD or random projections, we can efficiently leverage previously computed subspaces and the estimation error to better adjust the projection. We also incorporate subspace shifts into Adam’s first and second momentum update rules, ensuring proper alignment with coordinate changes using a Projection-Aware Optimizer (Robert et al., 2025). Additionally, we recover and scale (Recovery Scaling) the gradient information lost during low-rank projection to boost performance by utilizing the scaling information of the low-rank optimizer Chen et al. (2025); Zhu et al. (2025).
To summarize, SubTrack++ is a projection-aware geometry-based approach that supports full-parameter training and incorporates recovery scaling, offering superior time efficiency compared to SVD or PowerSGD methods (e.g., GaLore (Zhao et al., 2024), FiRA (Chen et al., 2025), LDAdam (Robert et al., 2025)), while maintaining GaLore’s memory footprint. It outperforms online PCA methods like Online Subspace Descent (Liang et al., 2024) in subspace tracking and achieves SOTA convergence and evaluation loss across all strong baselines; Comparison provided in Figure 1.
2 SubTrack++
Challenges of Low-Rank Optimization. Projecting gradients into a low-rank subspace reduces memory overhead and enables scalable LLM training, but it introduces important trade-offs. First, gradient subspaces can be unstable and require adaptive tracking; while SVD-based methods can capture these shifts (Zhao et al., 2024; Chen et al., 2025), they are computationally expensive. To address this, recent work has explored cheaper approximations and random projections (Robert et al., 2025; Liang et al., 2024; Zhu et al., 2025; He et al., 2025). Furthermore, optimizers like Adam assume a fixed coordinate system, so any changes in the subspace must be reflected in their internal states to preserve consistent momentum updates. Finally, low-rank projections inherently discard some gradient components, that recovering and utilizing these discarded signals can lead to improved performance (Chen et al., 2025; Zhu et al., 2025).
Overview. We propose SubTrack++ , a memory- and time-efficient method for full, simultaneous parameter tuning that embeds geometric insights into low-rank optimization. It improves efficiency and performance via three core components: 1) Subspace Tracking—Grassmannian-based updates that refine projections using estimation error and subspace history; 2) Projection-Aware Optimizer—adjusting Adam’s internal states to account for shifting subspaces; and 3) Recovery Scaling—rescaling discarded gradient components to reclaim lost information.
Subspace Tracking. In SubTrack++ , subspace estimation is framed as selecting a point on the Grassmannian—the manifold of all $d$-dimensional subspaces in an $n$-dimensional space (Bendokat et al., 2024)—which is the formal definition and goal of low-rank projections. This geometric perspective offers three key benefits: 1) By moving along Grassmann geodesics, it refines the subspace using prior subspace and estimation error, avoiding full reinitialization as done in (Zhao et al., 2024; Chen et al., 2025; He et al., 2025). 2) Updates rely on lightweight algebraic operations, minimizing time and memory costs. 3) Controlled subspace shifts improve robustness against noise and abrupt changes.
The initial subspace is computed using SVD as shown in (1). $G_{0}\in\mathbb{R}^{m\times n}$ is the gradient matrix at step 00; $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
$\small G_{0}=USV^{\top}\approx\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top}$ (1)
At each step, gradients are projected onto the subspace of left singular vectors if $m\leq n$, and right singular vectors otherwise; optimizing memory usage (Zhao et al., 2024). We assume $m\leq n$ without loss of generality, so the subspace is represented by $S_{t}\in\mathbb{R}^{m\times r}$ ($S_{0}=[u_{1},...,u_{r}]$), an orthonormal basis spanning the top-$r$ directions. The gradient is projected as $\widetilde{G}_{t}=S_{t}^{\top}G_{t}\in\mathbb{R}^{r\times n}$, and the optimizer operates in this reduced space, significantly lowering memory and state overhead.
To track subspace drift, we adjust the core subspace every $k$ steps, defined as the subspace update interval, by minimizing its Euclidean distance to the current gradient. The objective is formalized in the cost function (2).
$\small F(S_{t})=\min_{A}\|S_{t}A-G_{t}\|^{2}_{F},$ (2)
where $A$ is the solution to the least squares problem. The derivative of (2) with respect to $S_{t}$ is given in (3), and the residual $R={G_{t}}-S_{t}A$ lies in the orthogonal complement of $S_{t}$. To update the subspace, we calculate the tangent vector $\nabla F$ on the Grassmannian, as shown in (4) based on Edelman et al. (1998), where the second equality holds because $R$ is orthogonal to $S_{t}S_{t}^{\top}$.
$\small\frac{\partial F}{\partial S_{t}}=2(S_{t}A-G_{t})A^{\top}=-2RA^{\top}$ (3)
$\small\nabla F=(I-S_{t}S_{t}^{\top})\frac{\partial F}{\partial S_{t}}=\frac{\partial F}{\partial S_{t}}=-2RA^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ (4)
The tangent vector $\nabla F$ provides the direction for adjusting the subspace by accounting for the error lying in the orthogonal complement. However, to minimize changes to the subspace, SubTrack++ first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (4). This approximation is then used to update the subspace.
As shown by Edelman et al. (1998); Bendokat et al. (2024), from which we have included the necessary definitions and theorem in section 3, we can move along a Grassmannian geodesic guided by the computed tangent vector, taking a step of size $\eta$, as presented in (5).
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$ (5)
This update rule preserves the orthonormality of $S_{t+1}$, ensuring it remains on the Grassmannian. The last term in (4), projects the previous subspace onto the orthogonal complement of $\widehat{V}_{F}$, ensuring that the portion of $S_{t}$ which has not been updated in this step is still included.
Algorithm 1 SubTrack++
(Subspace Tracking, Projection-Aware Optimizer, Recovery Scaling, Regular Adam)
[…]
  $S_{0}\leftarrow U[:,:r]$ , where $U,S,V\leftarrow\text{SVD}(G_{0})$Initializing First Subspace}
  for $t=0,\ldots,T$ do
     if $t$ mod $k==0$ then
        $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$, and $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
        $M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$$\widetilde{G}_{t}=S_{t}^{\top}G_{t}$: low-rank projection of $G_{t}$} $\mathcal{V}_{t}\leftarrow\beta_{2}\cdot[(1-\beta_{2}^{t-1})|(S_{t}^{\top}S_{t-1})^{2}\cdot(\mathcal{V}_{t-1}-M_{t-1}^{2})+(S_{t}^{\top}S_{t-1}\cdot M_{t-1})^{2}|]+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$
     else
        $S_{t}=S_{t-1}$
        $M_{t}\leftarrow\beta_{1}\cdot M_{t-1}+(1-\beta_{1})\cdot\widetilde{G}_{t}$ $\mathcal{V}_{t}\leftarrow\beta_{2}\cdot\mathcal{V}_{t-1}+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$
     end if
      $\widetilde{G}_{t}^{O}=M_{t}\oslash\sqrt{\mathcal{V}_{t}+\epsilon}$, $\widehat{G}_{t}=S_{t}\widetilde{G}_{t}^{O}$$\widetilde{G}_{t}^{O}$: optimizer’s output, $\widehat{G}_{t}$: projected-back gradients }
[…]
     if $\frac{\Lambda_{t}}{\Lambda_{t-1}}>\zeta$ then $\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$
     $W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}$ $-\alpha\cdot\Lambda_{t}$
  end for
Projection-Aware Optimizer. In Adam, the first and second momentum update rule is shown in (6) and (7), respectively (reminder: $\widetilde{G}_{t}$ is the projection of gradient into low-rank subspace).
$\small M_{t}\leftarrow\beta_{1}\cdot M_{t-1}+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (6)
$\small\mathcal{V}_{t}\leftarrow\beta_{2}\cdot\mathcal{V}_{t-1}+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$ (7)
Inspired by methods such as LDAdam (Robert et al., 2025), we emphasize the importance of updating optimizer states in a projection-aware manner to account for shifting subspaces; otherwise, misaligned projections can distort the optimizer’s performance. To address this, at each subspace update step, we modify Adam’s original update rules in (6) and (7), replacing them with projection-aware counterparts represented in (8) and (9), which reflect subspace changes into optimizer statistics (Robert et al., 2025). Further details regarding this projections can be found in Appendix C.
$\small M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (8)
$\small\mathcal{V}_{t}\leftarrow\beta_{2}\cdot[(1-\beta_{2}^{t-1})|(S_{t}^{\top}S_{t-1})^{2}\cdot(\mathcal{V}_{t-1}-M_{t-1}^{2})+(S_{t}^{\top}S_{t-1}\cdot M_{t-1})^{2}|]+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$ (9)
These projection-aware update rules enables Adam optimizer to track optimization dynamics precisely as the subspace evolves, achieving significantly better practical performance.
[…]
$\small W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}-\alpha\cdot\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ (10)
Here $\alpha$ is the learning rate and $\phi_{t}(G_{t})$ is the scaling factor computed based on the low-rank gradient representation, $\widetilde{G}_{t}$, and the optimizer’s processed output, $\widetilde{G}_{t}^{O}$ as:
$\small\phi_{t}(G_{t})_{i}=\frac{\|\widetilde{G}_{t,:,i}^{O}\|}{\|\widetilde{G}_{t,:,i}\|}$ (11)
As suggested by Fira (Chen et al., 2025), this additional term must be controlled using a gradient-clipping-like method to ensure training stability. Building on their observation, we also limit the growth rate of $\Lambda_{t}=\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ by a factor $\zeta$, and apply a correction whenever it exceeds this threshold as:
$\small\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$ (12)
By incorporating a geometry-aware perspective into low-rank optimization and applying this approach across all components of the training pipeline, SubTrack++ demonstrates robust and stable training, and achieves state-of-the-art performance while maintaining minimal memory footprint and wall-time. The overall flow of operations in SubTrack++ is illustrated in Algorithm 1.
3 Theoretical Analysis
In this section, we analyze the convergence of Grassmannian Subspace Tracking applied in SubTrack++ using theoretical analysis. To begin, the general weights update rule is as follows:
$\small W_{t}=W_{0}-\alpha\cdot\sum_{t^{\prime}=0}^{t^{\prime}=t-1}\widehat{G}_{t^{\prime}}$ (13)
As previously mentioned, we use left projection if $m\leq n$, where $m$ and $n$ are the dimensions of the gradient matrix, and vice versa. Thus, $\widehat{G}_{t^{\prime}}$ can be computed as shown in (14).
$\small\widehat{G}_{t^{\prime}}=\begin{cases}S_{t^{\prime}}\rho_{t^{\prime}}(S_{t^{\prime}}^{\top}G_{t^{\prime}}),&\text{if $m\leq n$}\\ \rho_{t^{\prime}}(G_{t^{\prime}}S_{t^{\prime}})S_{t^{\prime}}^{\top},&\text{otherwise}\end{cases}$ (14)
Here, $S_{t^{\prime}}$ is the projection matrix that projects the gradient onto the subspace, and $\rho_{t^{\prime}}$ is representing the entry-wise regularizer used in the optimizer. If we use the full projection, then $\widehat{G}_{t^{\prime}}$ will be computed as shown in (15); where $S_{t^{\prime}}^{l}$ and $S_{t^{\prime}}^{r}$ are the rank-$r$ left and right projection matrices.
$\small\widehat{G}_{t^{\prime}}=S_{t^{\prime}}^{l}\rho_{t^{\prime}}({S_{t^{\prime}}^{l}}^{\top}G_{t^{\prime}}S_{t^{\prime}}^{r}){S_{t^{\prime}}^{r}}^{\top}$ (15)
Definition 3.1 (L-continuity).
A function $f(X)$ has Lipschitz-continuity (L-continuity) if for any $X_{1}$ and $X_{2}$, $\|f(X_{2})-f(X_{1})\|_{F}\leq L\|X_{2}-X_{1}\|_{F}$
{restatable}
[Convergence of Grassmannian Subspace Tracking]theoremconvergence Suppose gradient has the following form with functions $A_{i}$, $B_{i}$, and $C_{i}$ being L-continuous as per Def. 3.1 with constants $L_{A}$, $L_{B}$, and $L_{C}$ w.r.t. weight matrix $W_{t}$; and $\|W_{t}\|_{F}\leq M$; where $W_{t}$ denotes the weight matrix at step $t$, and $M$ is a scalar value,
$\small G=\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i}.$
Now, define $\widehat{B}_{i,t}=(S_{i,t}^{l})^{\top}B_{i}(W_{t})S_{i,t}^{l}$ and $\widehat{C}_{i,t}=(S_{i,t}^{r})^{\top}C_{i}(W_{t})S_{i,t}^{r}$, where $S_{i,t}^{l}$ and $S_{i,t}^{r}$ are the rank-$r$ left and right projection matrices; $B_{i}(W_{t})$ and $C_{i}(W_{t})$ denote the dependence of $B_{i}$ and $C_{i}$ on the weight matrices $W_{t}$. Further letting $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$, and $\kappa_{t}=\frac{1}{N}\sum_{i}\lambda_{min}(\widehat{B}_{i,t})\lambda_{min}(\widehat{C}_{i,t})$, where $\lambda_{min}(\cdot)$ denotes the minimum eigenvalue over each batch, and $N$ representing the number of samples in a batch. Assuming that the projection matrices remain constant during the training. Then for learning-rate $\mu$ and $min(\kappa_{t})>(L_{A}+2L_{B}L_{C}M^{2})$, subspace tracking, with $\rho_{t}\equiv 1$ (the element-wise regularizer of the optimizer) satisfies:
$\|P_{t}\|_{F}\leq[1-\mu(\kappa_{t-1}-L_{A}-2L_{B}L_{C}M^{2})]\|P_{t-1}\|_{F}.$
That is, $P_{t}\rightarrow 0$ and it converges.
The proof of Theorem 3 is provided in Appendix A, based on Zhao et al. (2024). While both GaLore and SubTrack++ assume the subspace remains unchanged for the proof of convergence, GaLore must limit these updates to ensure convergence, as each update can potentially change the entire subspace. In contrast, SubTrack++ leverages rank-$1$ updates to the subspace, preventing drastic changes with each update. While a deeper analysis of slowly changing subspaces and their impact on convergence remains an open problem, in practice, this allows SubTrack++ to perform more frequent updates.
Here we investigate the Grassmannian update rule presented in (5), which is a direct application of Grassmann geometry (Edelman et al., 1998; Bendokat et al., 2024).
Definition 3.2 (Exponential Map).
The exponential map $\exp_{p}:T_{p}M\to M$ on a Riemannian manifold $M$ is a mapping that assigns the point $\gamma(1)\in M$ to each tangent vector $\Delta\in T_{p}M$, where $T_{p}M$ is the tangent space of $M$ at $p$, and $\gamma$ is the unique geodesic originating at $p$ with initial velocity $\Delta$. This map establishes a relationship between geodesics and the Riemannian exponential, such that $\gamma(t)=\exp_{p}(t\Delta)$ for $t\in\mathbb{R}$.
Definition 3.3 (Stiefel Manifold).
The Stiefel manifold $St(n,p)$, parametrizes the set of all $n\times p$ orthonormal matrices $U$, each representing a rank-$p$ subspace of $\mathbb{R}^{n}$.
Definition 3.4 (Grassmann Manifold).
The Grassmannian manifold $Gr(n,p)$ parametrizes the set of all $p$-dimensional subspaces of $\mathbb{R}^{n}$. Each point can be represented by a projection matrix $P=UU^{\top}$, where $U\in St(n,p)$.
{restatable}
[Grassmann Exponential]theoremupdaterule Let $P=UU^{\top}\in\text{Gr}(n,p)$ be a point on the Grassmannian, where $U\in\text{St}(n,p)$ is the orthonormal basis of the corresponding subspace. Consider a tangent vector $\Delta\in T_{P}\text{Gr}(n,p)$, and let $\Delta_{U}^{\text{hor}}$ denote the horizontal lift of $\Delta$ to the horizontal space at $U$ in the Stiefel manifold $\text{St}(n,p)$. Suppose the thin SVD of $\Delta_{U}^{\text{hor}}$ is given by $\Delta_{U}^{\text{hor}}=\hat{Q}\Sigma V^{\top},$ where $\hat{Q}\in\text{St}(n,r)$, $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the nonzero singular values of $\Delta_{U}^{\text{hor}}$ with $r=\min(p,n-p)$, and $V\in\text{St}(p,r)$. The Grassmann exponential map, representing the geodesic emanating from $P$ in the direction $\Delta$, is given by:
$\small\text{Exp}_{P}^{\text{Gr}}(t\Delta)=\big{[}UV\cos(t\Sigma)V^{\top}+\hat{Q}\sin(t\Sigma)V^{\top}+UV_{\perp}V_{\perp}^{\top}\big{]},$
where $V_{\perp}\in\mathbb{R}^{p\times(p-r)}$ is any orthogonal complement of $V$.
The proof of Theorem 3 can be found in Appendix B. Leveraging this theorem and our notation in section 2, one can easily verify that the subspace update rule is as follows:
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
This update rule generally converges to a stable subspace if the step size $\eta$ decreases over time (Balzano et al., 2011). However, a decreasing step size can impair the ability to accurately track and adapt to subspace changes. Consequently, SubTrack++ uses a constant step size to effectively adjust subspaces. This approach does not hinder convergence, as proved in Theorem 3, which guarantees convergence as long as changes are controlled to maintain the stable subspace assumption.
4 Experiments and Results
To ensure fair comparison, we evaluated SubTrack++ across diverse models and datasets through pre-training and fine-tuning, measuring key metrics critical to LLM democratization.
Pre-Training Experiments. We pre-trained several Llama-based models on the C4 dataset, with results in Table 1. To ensure a fair comparison, we benchmarked against a diverse set of baselines. BAdam (Luo et al., 2024) is notably efficient in both memory and time, but its reliance on partial parameter updates leads to reduced performance. GaLore (Zhao et al., 2024) supports full-parameter training with low memory usage, though this comes at the cost of more than 150% increase in wall-time relative to AdamW in large models.
Table 1: We compare evaluation loss for pre-training Llama-based architectures on the C4 dataset over 10K iterations. SubTrack++ outperforms all other baselines in nearly every configuration. The best results are marked in bold, with the second-best performance underlined. LDAdam could not be run on the 7B configuration due to an out-of-memory error with our available resources.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 3.41 3.25 3.40 4.61 4.52 4.30
GaLore (Zhao et al., 2024) 4.02 3.61 3.62 6.53 6.57 5.55
BAdam (Luo et al., 2024) 7.86 7.08 7.62 7.28 7.12 6.76
Online Subspace Descent (Liang et al., 2024) 4.18 3.88 4.09 6.79 6.85 5.69
LDAdam (Robert et al., 2025) 3.52 3.44 3.67 4.70 4.39 OOM
Fira (Chen et al., 2025) 3.80 3.55 3.56 6.31 6.50 6.83
SubTrack++ (Ours) 3.43 3.24 3.29 4.51 4.49 4.62
As another relative baseline, Online Subspace Descent (Liang et al., 2024) uses Online PCA to track gradient subspace. LDAdam (Robert et al., 2025) and Fira (Chen et al., 2025) deliver strong performance at the cost of increasing wall-time even more than GaLore. Across these methods, SubTrack++ achieves SOTA results while using 14% less memory and running 43% faster than LDAdam, the most accurate baseline on 1B model. Hyperparameter details are provided in Appendix G, with a detailed wall-time and memory analysis in Appendix F.
Figure 3: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ . While Grassmannian tracking alone matches GaLore’s step-wise convergence, it significantly reduces wall-time, as shown in Figure LABEL:fig:modular.
Figure LABEL:fig:modular compares pre-training performance of a 1B-parameter Llama model across several baselines. SubTrack++ converges fastest in both steps and wall-time, emphasizing the power of geometry-aware optimization in boosting performance, reducing resource demands, and democratizing access to SOTA training methods.
Ablation Study. We conducted an ablation study to evaluate the individual and combined effects of projection-aware optimizers and recovery scaling, integrated with Grassmannian subspace tracking on the 1B Llama model. We use the experimental settings outlined in Table 10. As shown in Figure 3, projection-aware optimizer and recovery scaling independently yield substantial improvements over pure Grassmannian subspace tracking, reducing loss from 6.53 to 5.43 and 5.28, respectively; while their combination, SubTrack++ , achieves 4.51 surpassing all baselines. Notably, these gains are achieved with minimal increases in runtime and memory.
Fine-Tuning Experiments. RoBERTa-Base and RoBERTa-Large are fine-tuned on GLUE and SuperGLUE tasks; with the results presented in Table 4 and 5, respectively. More details and hyperparameters are provided in Appendix E.
Time and Space Complexity. Table 2 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$. GaLore (Zhao et al., 2024) and Fira (Chen et al., 2025) periodically perform SVD to estimate the underlying subspace, while LDAdam (Robert et al., 2025) relies on the faster PowerSGD to update the subspace at every iteration. In contrast, SubTrack++ employs Grassmannian-based subspace tracking at the same frequency as GaLore and Fira. Comparing the time complexities of these methods, highlights why SubTrack++ is significantly more efficient than SVD-based methods. A breakdown of the subspace update time complexity for SubTrack++ is shown in Appendix D. Additionally, the memory required for storing optimizer states in SubTrack++ , is equivalent to GaLore and other baselines.
Table 2: The optimizer’s state parameter count and subspace update time complexity across baselines, given a gradient matrix of dimension $m\times n$ and projection rank $r$ where $r\ll m\leq n$. LDAdam updates the subspace at every iteration, while other methods update it every $k$ steps.
Optimizer Mem. Subspace Update Time
Adam $2mn$-
LDAdam $mr+2nr$ $O(mnr)$
GaLore, Fira $mr+2nr$ $O(nm^{2})$
SubTrack++ $mr+2nr$ $O(mnr)$
(a) Ours, SF=1
(b) GaLore, SF=1
(c) Ours, SF=3
(d) GaLore, SF=3
Figure 5: Comparison of Grassmannian subspace tracking (Ours) (a, c) and GaLore’s SVD (b, d) on the Ackley Function over 100 optimization steps, with a subspace update interval of 10. SF stands for scale factor; with a scale factor of 1, GaLore fails to reach the global minimum due to abrupt jumps. At a scale factor of 3, while the minimum is reached, the jump length increases. This demonstrates SVD’s sensitivity to noise and abrupt changes, highlighting the robustness of our Subspace Tracking method with its controlled subspace updates.
Robust Subspace Tracking. Relying on SVD for subspace updates makes methods sensitive to noise and abrupt changes (He et al., 2025). Figure 5 compares Grassmannian subspace tracking with GaLore’s SVD on the Ackley function, highlighting how SVD causes erratic jumps, while our subspace tracking ensures robust optimization. GaLore struggles to reach the global minimum within 100 steps at scale factor 1, and although increasing the scale factor to 3 improves performance, it amplifies jumps that hinder convergence in non-convex settings, revealing sensitivity to hyperparameters, noise, and abrupt changes.
5 Related Works
Parameter-Efficient Training. Several works aim to improve the efficiency of training LLMs, addressing a growing demand as their popularity rapidly increases. Popular LoRA (Hu et al., 2021) significantly reduces memory requirements for fine-tuning LLMs by leveraging two low-rank trainable low-rank matrices. Dettmers et al. (2024) employ quantization techniques and paged optimizers to further reduce memory usage. Additionally, Yaras et al. (2024) introduce Deep LoRA to address overfitting issues and reducing the need for precise tuning of the rank parameter. Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024). Miles et al. (2024) propose compressing intermediate activations and reconstructing them during backpropagation to enhance memory efficiency. Additionally, Hao et al. (2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient. BAdam (Luo et al., 2024) leverages the block coordinate descent framework to reduce memory consumption while maintaining capabilities comparable to Adam.
Gradient Low-Rank Projection. Several approaches aim to reduce optimizer states, as optimizers like Adam (Kingma and Ba, 2017) account for a significant portion of memory footprint (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022). MicroAdam (Modoranu et al., 2024) tackles this by compressing the gradient space and utilizing the compression error through feedback loops. Adam-mini (Zhang et al., 2024) partitions model into blocks, assigning a single learning rate to each block to preserve performance while saving memory. Gur-Ari et al. (2018); Schneider et al. (2024); Yaras et al. (2023) suggest that a substantial portion of gradients lies within a largely consistent subspace. GaLore (Zhao et al., 2024) leverages this fact to reduce the optimizer’s memory by projecting gradients into a low-rank subspace and then projecting them back for full parameter tuning. This approach has been integrated with other methods regarding efficient LLM training (Li et al., 2024; Rajabi and Rambhatla, 2024b). However, not all layers’ gradients evolve within a stable low-rank subspace. Jaiswal et al. (2024); Rajabi and Rambhatla (2024d) identify layers where gradients evolve within a low-dimensional subspace and fine-tune only those layers, freezing the others to avoid inefficient low-rank updates. Grass (Muhamed et al., 2024) reduces memory usage by applying sparse projection matrices to the gradient. Ramesh et al. (2024) dynamically select and update a subset of parameters, for a fast and memory-efficient training. FIRA (Chen et al., 2025) utilize a norm-based scaling method along with GaLore to maintain performance comparable to full-rank training. GoLore (He et al., 2025) addresses GaLore’s convergence issues and employ random projection in latter steps as a solutions. LDAdam (Robert et al., 2025) performs optimization within lower-dimensional subspaces, incorporating a projection-aware optimization update rule and a generalized error feedback mechanism. APOLLO (Zhu et al., 2025) approximates channel-wise learning rate scaling based on random projection. Also, Liang et al. (2024) introduce a dynamically evolving projection matrix updated via online PCA, enhancing the model’s ability to navigate the parameter space efficiently without relying on expensive SVD, a similar approach is also proposed in Rajabi and Rambhatla (2024a).
Grassmannian and Subspace Tracking. A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve. Balzano et al. (2011); Rajabi and Rambhatla (2024c) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed. Zhang and Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces. Furthermore, Blocker et al. (2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively.
6 Discussion and Conclusion
We propose SubTrack++ , a time- and memory-efficient approach that projects gradients into a low-rank subspace and uses Grassmannian subspace tracking to preserve the computed subspace while incorporating gradient components from the orthogonal complement. By integrating projection-aware optimizers that reflect subspace changes in Adam’s internal statistics and utilizing the gradient information lost during low-rank projection, SubTrack++ achieves state-of-the-art convergence and accuracy across all baselines. While Grassmannian subspace tracking integrates seamlessly with various optimizers as a plug-and-play module, extending projection-aware optimization beyond the Adam family requires further design and investigation. Also benchmarking on models with more than 7B parameters was not feasible regarding our limited time and resources.
Appendix A Convergence of SubTrack++
[…]
proof. To demonstrate that SubTrack++ converges to the global minimum during training, we begin by deriving the recursive form of the gradients.
Let $\otimes$ denote the Kronecker product. Then, $vec(AXB)=(B^{\top}\otimes A)vec(X)$.
By applying $vec$ to the gradient form given in the theorem, we obtain:
$g_{t}=vec(G_{t})=vec(\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i})=a_{t}-D_{t}w_{t}$ (16)
where $g_{t}:=vec(G_{t})$, $w_{t}:=vec(W_{t})$, $a_{t}:=\frac{1}{N}\sum_{i}vec(A_{i,t})$, and $D_{t}=\frac{1}{N}\sum_{i}C_{i,t}\otimes B_{i,t}$.
As defined in the theorem, let $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$. Its vectorized form can be expressed using the Kronecker product as follows:
$\begin{gathered}p_{t}=vec(P_{t})=vec({S_{t}^{l}}^{\top}G_{t}S_{t}^{r})=({S_{t}^{r}}^{\top}\otimes{S_{t}^{l}}^{\top})vec(G_{t})\\ \quad={(S_{t}^{r}\otimes S_{t}^{l})}^{\top}vec(G_{t})={(S_{t}^{r}\otimes S_{t}^{l})}^{\top}g_{t}\end{gathered}$ (17)
Now recalling $\widehat{G}_{t}$ from (15), it can be written as:
$\widehat{G}_{t}=S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}S_{t}^{r}{S_{t}^{r}}^{\top}$
Thus, its vectorized form will be:
$\begin{gathered}vec(\widehat{G}_{t})=\widehat{g}_{t}=vec(S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}S_{t}^{r}{S_{t}^{r}}^{\top})=vec(S_{t}^{l}P_{t}{S_{t}^{r}}^{\top})\\ \quad=(S_{t}^{r}\otimes S_{t}^{l})vec(P_{t})=(S_{t}^{r}\otimes S_{t}^{l})p_{t}\end{gathered}$ (18)
This is where the constant subspace assumption becomes necessary. To derive the recursive form of $g_{t}$, we assume that the projection matrices remain fixed throughout training, i.e., $S_{t}^{r}=S^{r}$ and $S_{t}^{l}=S^{l}$. Consequently, we can restate equations (17) and (18) as follows:
$p_{t}={(S^{r}\otimes S^{l})}^{\top}g_{t}$ (19)
$\widehat{g}_{t}=(S^{r}\otimes S^{l})p_{t}$ (20)
Then we can write the recursive form of $g_{t}$:
$\begin{gathered}g_{t}=a_{t}-D_{t}w_{t}=(a_{t}-a_{t-1})+(D_{t-1}-D_{t})w_{t}+a_{t-1}-D_{t-1}w_{t}\\ \quad=e_{t}+a_{t-1}-D_{t-1}(w_{t-1}+\mu\widehat{g}_{t-1})=e_{t}+g_{t-1}-\mu D_{t-1}\widehat{g}_{t-1}\end{gathered}$ (21)
where $e_{t}:=(a_{t}-a_{t-1})+(D_{t-1}-D_{t})w_{t}$.
Note that in deriving (21), we utilized the general form of the weight update rule, $w_{t+1}=w_{t}-\mu g_{t}$, which can be rewritten as $w_{t}=w_{t+1}+\mu g_{t}$. By applying this rule along with (16), we arrive at the second equality in (21) as follows:
$\begin{gathered}g_{t}=a_{t}-D_{t}w_{t}=a_{t}-D_{t}w_{t}-g_{t-1}+g_{t-1}\\ \quad=a_{t}-D_{t}w_{t}-a_{t-1}+D_{t-1}w_{t-1}+a_{t-1}-D_{t-1}w_{t-1}\\ \quad=a_{t}-D_{t}w_{t}-a_{t-1}+D_{t-1}(w_{t}+\mu g_{t-1})+a_{t-1}-D_{t-1}(w_{t}+\mu g_{t-1})\\ \quad=a_{t}-D_{t}w_{t}-a_{t-1}+D_{t-1}w_{t}+\mu D_{t-1}g_{t-1}+a_{t-1}-D_{t-1}w_{t}-\mu D_{t-1}g_{t-1}\\ \quad=a_{t}-a_{t-1}+(D_{t-1}-D_{t})w_{t}+a_{t-1}-D_{t-1}\end{gathered}$
To obtain $p_{t}$ from this recursive formulation, we can left-multiply by ${(S^{r}\otimes S^{l})}^{\top}$, as shown in (20):
$\begin{split}p_{t}={(S^{r}\otimes S^{l})}^{\top}e_{t}+{(S^{r}\otimes S^{l})}^{\top}g_{t-1}-\mu{(S^{r}\otimes S^{l})}^{\top}D_{t-1}\widehat{g}_{t-1}\end{split}$ (22)
Now, based on (19) and (20), $p_{t}$ can be written as:
$p_{t}={(S^{r}\otimes S^{l})}^{\top}e_{t}+p_{t-1}-\mu{(S^{r}\otimes S^{l})}^{\top}D_{t-1}{(S^{r}\otimes S^{l})}p_{t-1}$ (23)
Let define:
$\begin{gathered}\widehat{D}_{t}:={(S^{r}\otimes S^{l})}^{\top}D_{t}(S^{r}\otimes S^{l})=\frac{1}{N}\sum_{i}{(S^{r}\otimes S^{l})}^{\top}(C_{i,t}\otimes B_{i,t})(S^{r}\otimes S^{l})\\ \quad=\frac{1}{N}\sum_{i}({S^{r}}^{\top}C_{i,t}S^{r})\otimes({S^{l}}^{\top}B_{i,t}S^{l})\end{gathered}$ (24)
Then we can expand (23) and show that:
$p_{t}=(I-\mu\hat{D}_{t-1})p_{t-1}+(S^{r}\otimes S^{l})^{\top}e_{t}$ (25)
Note that $S^{l}$ and $S^{r}$ are orthonormal matrices. This is ensured because the subspace is initialized using the SVD of $G_{0}$, and the Grassmannian update rule provided in (5) preserves the orthonormality of the subspace matrices throughout training. Since $S^{l}$ and $S^{r}$ are orthonormal, we have ${S^{l}}^{\top}S^{l}=I$ and ${S^{r}}^{\top}S^{r}=I$. Consequently, we can bound the norm of the second term in (25) as follows:
$\|(S^{r}\otimes S^{l})^{\top}e_{t}\|_{2}=\|vec({S^{l}}^{\top}E_{t}S^{r})\|_{2}=\|{S^{l}}^{\top}E_{t}S^{r}\|_{F}\leq\|E_{t}\|_{F}$ (26)
Here $E_{t}$ is the matrix form of $e_{t}$, and as declared before, $e_{t}:=(a_{t}-a_{t-1})+(D_{t-1}-D_{t})w_{t}$, thus:
$E_{t}:=\frac{1}{N}\sum_{i}(A_{i,t}-A_{i,t-1})+\frac{1}{N}\sum_{i}(B_{i,t-1}W_{t}C_{i,t-1}-B_{i,t}W_{t}C_{i,t})$ (27)
Next, we need to find an upper bound for the norm of each term in (27) to establish an upper bound for $\|E_{t}\|_{F}$. Based on the assumptions of the theorem, $A_{i}$, $B_{i}$, and $C_{i}$ exhibit L-Lipschitz continuity with constants $L_{A}$, $L_{B}$, and $L_{C}$, respectively. Additionally, $\|W_{t}\|_{F}$ is bounded by a scalar $M$. We have:
$\|A_{t}-A_{t-1}\|_{F}$$\leq$ $L_{A}\|W_{t}-W_{t-1}\|_{F}=\mu L_{A}\|\tilde{G}_{t-1}\|_{F}\leq\mu L_{A}\|P_{t-1}\|_{F}$ (28)
In the first equality, we apply (13), while the last equality holds due to (20) and the orthonormality of the projection matrices. The subsequent two inequalities can be derived similarly using these equations.
$\begin{gathered}\|(B_{t}-B_{t-1})W_{t}C_{t-1}\|_{F}\leq L_{B}\|W_{t}-W_{t-1}\|_{F}\|W_{t}\|_{F}\|C_{t-1}\|_{F}\\ \quad=\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}\end{gathered}$ (29)
$\begin{gathered}\|B_{t}W_{t}(C_{t-1}-C_{t})\|_{F}\leq L_{C}\|B_{t}\|_{F}\|W_{t}\|_{F}\|W_{t-1}-W_{t}\|_{F}\\ \quad=\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}\end{gathered}$ (30)
We can now derive the bound for $\|E_{t}\|_{F}$ as follows:
$\begin{gathered}\|E_{t}\|_{F}\leq\mu L_{A}\|\tilde{G}_{t-1}\|_{F}\leq\mu L_{A}\|P_{t-1}\|_{F}+\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}+\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}\\ \quad=\mu(L_{A}+2L_{B}L_{C}M^{2})\|P_{t-1}\|_{F}\end{gathered}$ (31)
To calculate the norm bound for the first term in (25), we first need to establish the bounds for $\widehat{D}_{t}$. This involves estimating the minimum eigenvalue of $\widehat{D}_{t}$.
[…]
this actually provides a lower bound for eigenvalues of $\widehat{D}_{t}$, thus:
$\lambda_{\max}(I-\mu\widehat{D}_{t-1})\leq 1-\frac{\mu}{N}\sum_{i}\gamma_{min,i,t-1}$ (33)
considering the definition of $\kappa_{t}$ in the theorem, we can now easily show that:
$\|P_{t}\|_{F}\leq[1-\mu(\kappa_{t-1}-L_{A}-2L_{B}L_{C}M^{2})]\|P_{t-1}\|_{F}.$
and completing the proof.
While SubTrack++ utilizes right/left projections to reduce memory consumption, the proof is presented using both projection matrices to ensure generality. Here, we demonstrate how the proof proceeds under the assumption $m\leq n$ (without loss of generality), which allows the use of the left projection matrix.
Using the left projection matrix, the current formulation of $P_{t}$, defined as $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$, simplifies to $P_{t}={S_{t}^{l}}^{\top}G_{t}$. Similarly, $\widehat{G}_{t}=S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}S_{t}^{r}{S_{t}^{r}}^{\top}$ reduces to $\widehat{G}_{t}=S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}$. From this point, the proof continues by substituting $S_{t}^{r}$ with the identity matrix, allowing the derivation of the vectorized forms of $g_{t}$, $\widehat{g}_{t}$, $p_{t}$, and related terms.
The remainder of the proof remains largely unaffected. It can be readily verified that the recursive formulation of $g_{t}$ is unchanged. Although the definition of $P_{t}$ is modified, it continues to satisfy the bounds required for convergence, ensuring that $P_{t}$ converges to 0 when the left projection matrix is used.
Appendix B Grassmann Exponential
[…]
proof. Using Grassmannina mathematics, we know that every $\Delta\in T_{P}Gr(n,p)$ is of the form
$\Delta=Q\begin{pmatrix}0&B^{\top}\\ B&0\end{pmatrix}Q^{\top}=\left[Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}Q^{\top},P\right]$ (34)
Then the lift of $\Delta\in T_{P}Gr(n,p)$ to $Q=\begin{pmatrix}U&U_{\perp}\end{pmatrix}$ can also be calculated explicitly as follows:
$\Delta^{\text{hor}}_{Q}=[\Delta,P]Q=Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}$ (35)
To resume our proof, we need to define the orthogonal group and specifying its tangent space.
Definition B.1 (Orthogonal Group).
The orthogonal group $O(n)$ is defined as the set of all $n\times n$ matrices $Q$ over $\mathbb{R}$ such that $Q^{\top}Q=QQ^{\top}=I_{n}$, where $Q^{\top}$ is the transpose of $Q$ and $I_{n}$ is the $n\times n$ identity matrix:
$O(n)=\{Q\in\mathbb{R}^{n\times n}\mid Q^{\top}Q=I_{n}=QQ^{\top}\}.$
Then the tangent space of the orthogonal group $O(n)$ at a point $Q$, denoted $T_{Q}O(n)$, is defined as the set of matrices of the form $Q\Omega$, where $\Omega\in\mathbb{R}^{n\times n}$ is a skew-symmetric matrix, i.e., $\Omega^{\top}=-\Omega$:
$T_{Q}O(n)=\{Q\Omega\mid\Omega\in\mathbb{R}^{n\times n},\Omega^{\top}=-\Omega\}.$
The geodesic from $Q\in O(n)$ in direction $Q\Omega\in T_{Q}O(n)$ is calculated via
$\text{Exp}_{Q}^{O}(tQ\Omega)=Q\exp_{m}(t\Omega),$ (36)
If $P\in Gr(n,p)$ and $\Delta\in T_{P}Gr(n,p)$ with $\Delta^{\text{hor}}_{Q}=Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}$, the geodesic in the Grassmannian is therefore
$\text{Exp}^{Gr}_{P}(t\Delta)=\pi^{OG}\left(Q\exp_{m}\left(t\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}\right)\right).$ (37)
where $\pi^{OG}$ is the projection from $O(n)$ to $Gr(n,p)$. If the thin SVD of $B$ is given by
$B=U_{\perp}^{\top}\Delta^{\text{hor}}_{U}=U_{\perp}^{\top}\hat{Q}\Sigma V^{\top}$
with $W:=U_{\perp}^{\top}\hat{Q}\in St(n-p,r),\Sigma\in\mathbb{R}^{r\times r},V\in St(p,r)$. Let $W_{\perp},V_{\perp}$ be suitable orthogonal completions. Then,
$\exp_{m}\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}=\begin{pmatrix}V&V_{\perp}&0&0\\ 0&0&W&W_{\perp}\end{pmatrix}\begin{pmatrix}\cos(\Sigma)&0&-\sin(\Sigma)&0\\ 0&I_{p-r}&0&0\\ \sin(\Sigma)&0&\cos(\Sigma)&0\\ 0&0&0&I_{n-p-r}\end{pmatrix}\begin{pmatrix}V^{\top}&0\\ V_{\perp}^{\top}&0\\ 0&W^{\top}\\ 0&W_{\perp}^{\top}\end{pmatrix},$
which leads to the desired result when inserted into (37). For more mathematical details, you can refer to Edelman et al. [1998], Bendokat et al. [2024], or other useful resources on Grassmann geometry.
Appendix C Projection-Aware Optimizer
When projecting into a lower-dimensional space and tracking coordinate changes, we typically use orthonormal projection matrices to represent the subspaces and their multiplications to effect a change of basis. While this works well for purely linear operations, Adam’s updates also incorporate non-linear elements.
Suppose the subspace changes at step $t$, transitioning from the subspace spanned by the orthonormal matrix $S_{t-1}$ to that spanned by $S_{t}$. Since both matrices are orthonormal—preserved by the Grassmannian update rule in (5)—the matrix $S_{t}^{\top}S_{t-1}$ represents the change of basis between the two subspaces. In other words, if $\mathbb{E}_{t-1}=(e_{t-1}^{1},\dots,e_{t-1}^{r})$ and $\mathbb{E}_{t}=(e_{t}^{1},\dots,e_{t}^{r})$ are orthonormal bases for the subspaces at steps $t-1$ and $t$, respectively, then the $i$th column of matrix $X$ transforms under the change of basis as $X^{i}_{t}=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle X^{j}_{t-1}$, where $X^{j}_{t-1}$ is the $j$th column of $X$ based on the basis of the time step $t-1$.
$\mathbf{E}_{t,\beta}\left[.\right]$ denotes the exponential time-weighted expectation at time $t$ with decay rate $\beta$. Following the reinterpretation by Robert et al. [2025], Adam’s first and second moment estimates can be expressed as $\widetilde{M}_{t}=\mathbf{E}_{t,\beta_{1}}\left[\widetilde{G}_{t}\right]$ and $\widetilde{\mathcal{V}}_{t}=\mathbf{E}_{t,\beta_{2}}\left[(\widetilde{G}_{t})^{2}\right]$, where $\widetilde{G}_{t}$ denotes the low-rank representation of the gradient at time step $t$. As shown in (38), the first moment estimate can be transformed under a change of basis using the change-of-basis matrix, $S_{t}^{\top}S_{t-1}$. Notably, $\langle\widetilde{G}_{t},e^{i}_{t}\rangle$ gives the $i$th column of $\widetilde{G}_{t}$ when the subspace has the basis $\mathbb{E}_{t}$. We use the superscripts to indicate a column of a matrix.
$\small\mathbf{E}_{t,\beta_{1}}\left[\langle\widetilde{G}_{t},e^{i}_{t}\rangle\right]=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle\mathbf{E}_{t,\beta_{1}}\left[\langle\widetilde{G}_{t},e^{j}_{t-1}\rangle\right]=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle\widetilde{M}_{t}^{j}=\left(S_{t}^{\top}S_{t-1}\widetilde{M}_{t}\right)^{i}$ (38)
Following the same approach, we can change the basis for the second moment estimate as described in (39).
$\small\mathbf{E}_{t,\beta_{2}}\left[\left(\langle\widetilde{G}_{t},e^{i}_{t}\rangle\right)^{2}\right]$ $=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\,\mathbf{E}_{t,\beta_{2}}\left[\left(\langle\widetilde{G}_{t},e^{j}_{t-1}\rangle\right)^{2}\right]$ (39)
$+\sum_{k\neq l}^{r}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\,\mathbf{E}_{t,\beta_{2}}\left[\langle\widetilde{G}_{t},e^{k}_{t-1}\rangle\langle\widetilde{G}_{t},e^{l}_{t-1}\rangle\right]$
$=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\widetilde{\mathcal{V}}_{t}^{j}$
$+\sum_{k\neq l}^{r}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}.$
In transitioning from the first equality to the second, we assume independence among the gradient coordinates. This enables us to approximate the covariance using a product of first-order moment estimates. This assumption is often reasonable in practice because we compute the SVD of the gradient and maintain an orthonormal subspace projection matrix, updating it along the Grassmannian geodesic to track the optimal subspace. Since SVD tends to diagonalize the covariance, the off-diagonal entries are typically negligible. Additionally, we clip any negative values to zero to ensure valid (non-negative) variance estimates. Moreover, to rewrite the second term in the final equality of (39), we employ the following equation:
$\sum_{k}\sum_{l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}$ (40)
$=\sum_{k}\langle e^{i}_{t},e^{k}_{t-1}\rangle^{2}\left(\widetilde{M}_{t}^{k}\right)^{2}\;+\;\sum_{k\neq l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}$
Given these, we can rewrite (39) as follows:
$\mathbf{E}_{t,\beta_{2}}\left[\left\langle\widetilde{G}_{t},e_{t}^{i}\right\rangle^{2}\right]$ $=\sum_{j}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\widetilde{\mathcal{V}}_{t}^{j}+$ (41)
$\left[\sum_{k}\sum_{l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}-\sum_{k}\langle e^{i}_{t},e^{k}_{t-1}\rangle^{2}\left(\widetilde{M}_{t}^{k}\right)^{2}\right]$
$=\sum_{j}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\left[\widetilde{\mathcal{V}}_{t}^{j}-\left(\widetilde{M}_{t}^{j}\right)^{2}\right]+\left(\langle e^{i}_{t},e^{j}_{t-1}\rangle\widetilde{M}_{t}{}\right)^{2}$
$=\left(\left(S_{t}^{\top}S_{t-1}\right)^{2}\left[\widetilde{\mathcal{V}}_{t}-\widetilde{M}_{t}^{2}\right]\right)^{i}+\left(\left(S_{t}^{\top}S_{t-1}\widetilde{M}_{t}\right)^{2}\right)^{i}$
By applying (38) and (41), we can directly derive the update rules for the projection-aware optimizer, as expressed in (8) and (9).
Appendix D Time Complexity Analysis
Table 3 presents the time complexity breakdown for the subspace update step in the SubTrack++ algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$. As outlined in Algorithm 1, the subspace update step begins by solving the least squares problem (2) to estimate the optimal update for $S_{t}$, the $m\times r$ orthonormal matrix. This operation has a time complexity of $O(mr^{2})$. Computing the residual and the partial derivative with respect to $S_{t}$ requires $O(mrn)$ and $O(mnr)$ time respectively. This is because the solution to the least squares problem, $A$, has shape $r\times n$ which is multiplied by $S_{t}$ in the residual $R=G_{t}-S_{t}A$, resulting in time complexity $O(mrn)$. The following operation for the partial derivative is $-2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity. The tangent vector computation (4) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$.
Table 3: Time Complexity for SubTrack++ Subspace Update
Computation Step Time
Cost function $O(mr^{2})$
Residual $O(mrn)$
Partial derivative $O(mnr)$
Tangent vector $\Delta F$ $O(m^{2}r)$
Rank-$1$ approximation of $\Delta F$ $O(mr^{2})$
Update rule $O(mr^{2})$
Overall $O(mnr)$
The rank-1 approximation step uses largest singular value from the SVD of the $m\times r$ tangent vector, and has time complexity of $O(mr^{2})$. Finally, the update rule as shown in (13) which has a time complexity of $O(mr^{2})$. The overall complexity of the algorithm is dominated by the matrix multiplication calculations of time complexity $O(mnr)$. However, unlike GaLore, since we avoid computing SVD operation on the $m\times n$ gradient matrix, which has complexity of $O(nm^{2})$, the overall update step in SubTrack++ is still more efficient with respect to time complexity.
Appendix E Fine-Tuning Experiments
As described, we examined SubTrack++ on fine-tuning RoBERTa-base and RoBERTa-large models to evaluate them on GLUE and SuperGLUE benchmarks. The results for GLUE task is summarized in Table 4, and SuperGLUE in 5.
Table 4: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for $r=8$. The performance is measured via Accuracy for SST-2 and RTE tasks, F1 for MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA. The best results are marked in bold, with the second-best performance underlined.
COLA STS-B MRPC RTE SST-2
Full-Rank 62.57 91.03 91.32 77.98 94.27
BAdam [Luo et al., 2024] 54.44 89.01 91.35 68.59 94.15
GaLore [Zhao et al., 2024] 58.54 90.61 91.30 74.37 94.50
LDAdam [Robert et al., 2025] 58.81 90.90 92.22 76.53 94.27
SubTrack++ (Ours) 58.29 90.90 91.84 76.53 90.83
Table 5: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with $r=8$. The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, and F1 for CB. The best results are marked in bold, with the second-best performance underlined.
BoolQ CB COPA WIC WSC AXg
Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30
GaLore [Zhao et al., 2024] 85.44 88.85 80.00 71.47 63.46 100.00
BAdam [Luo et al., 2024] 82.51 53.28 59.00 70.38 60.58 51.85
LDAdam Robert et al. [2025] 85.75 56.16 80.00 71.00 64.42 70.37
SubTrack++ (Ours) 84.68 89.66 88.00 70.06 63.46 100.00
Table 6: Hyperparameters of fine-tuning RoBERTa-Base on GLUE tasks.
SST-2 MRPC CoLA RTE STS-B
Shared Parameters Batch Size 16 16 32 16 16
# Epochs 30
Max Seq. Len. 512
Low-Rank Optimizer Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
Methods Parameters SubTrack Step-Size 0.01 3.0 5.0 15.0 10.0
Subspace Update Interval 500
Rank Config 8
$\alpha$ 2
BAdam Parameters Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
Block Switch Interval 100
Switch Mode Random
Table 7: Hyperparameters of fine-tuning RoBERTa-Large on SuperGLUE tasks.
BoolQ CB COPA WIC WSC AXg
Shared Parameters Batch Size 16
# Epochs 30
Learning Rate 1e-5
Max Seq. Len. 512
Low-Rank Optimizer SubTrack++ Step-Size 0.1 1.0 50.0 100.0 1.0 1.0
Methods Parameters Subspace Update Interval 500 100 100 500 250 100
Rank Config. 8
$\alpha$ 4
BAdam Parameters Block Switch Interval 100 50 50 100 50 50
Switch Mode Random
The hyperparameters for fine-tuning RoBERTa-base are detailed in Table 6, matching those reported in the GaLore [Zhao et al., 2024] for rank-$8$ subspaces, with a subspace update interval set at 500 iterations. We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al. [2024], as detailed in Table 7, with the exception that we fine-tuned each task for 30 epochs.
Appendix F Memory and Time Comparison
Table 8 presents the the peak memory consumption measured to compare SubTrack++ and other baselines. It shows that SubTrack++ requires on-par or better memory compared to GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025] while getting state-of-the-art results. As detailed in Table 2, all baselines except BAdam [Luo et al., 2024] use the same number of optimizer parameters; therefore, any differences in their peak memory consumption stem from variations in their runtime parameters.
Table 8: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset. The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with . SubTrack++ demonstrates better or on-par memory compared to other low-rank methods that allows full-parameter training.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 16.86 25.32 28.67 18.83 34.92 50.50
BAdam [Luo et al., 2024] 13.34 20.01 16.45 9.18 14.75 22.35
GaLore [Zhao et al., 2024] 16.89 25.52 27.85 14.74 26.03 36.00
Online Subspace Descent Liang et al. [2024] 16.61 25.86 28.76 18.45 32.57 33.10
LDAdam [Robert et al., 2025] 17.18 25.94 28.03 18.45 32.57 OOM
Fira [Chen et al., 2025] 16.39 24.99 27.33 14.68 25.62 47.84
SubTrack++ (Ours) 16.40 25.06 27.42 15.22 25.54 49.82
Additionally, Table 9 presents the wall-time consumed by each baseline across all model architectures during pre-training. Each run is configured to include exactly 10 subspace updates for SubTrack++ and other baselines employing periodic subspace updates. Specifically, models ranging from 60M to 3B use a subspace update interval of 200, resulting in 2,000 total iterations, while the 7B models use an interval of 500, yielding 5,000 iterations. Experiments for the 60M to 3B models are conducted on an NVIDIA A100 GPU, while the 7B model experiments are run on an NVIDIA RTX A6000.
Table 9: Wall-time comparison for pre-training LLaMA-based architectures of varying sizes on the C4 dataset. The number of iterations is set to ensure 10 subspace updates for methods using periodic subspace updates: 2k for models ranging from 60M to 3B, and 5k for the 7B model. SubTrack++ achieves the lowest wall-time among all baselines that support full-parameter training on large models. The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with . Since this can impact wall-time comparisons, it is more appropriate to compare runs within the same cluster.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 12726.2
BAdam [Luo et al., 2024] 511.3 779.2 961.6 798.6 1004.1 7283.7
GaLore [Zhao et al., 2024] 547.8 1094.2 1589.0 1729.5 2715.5 21590.4
Online Subspace Descent [Liang et al., 2024] 662.8 1228.2 1818.6 1438.7 1676.9 21590.4
LDAdam [Robert et al., 2025] 639.9 1342.2 2083.4 2780.9 4625.4 OOM
Fira [Chen et al., 2025] 635.3 1180.8 1729.7 1938.5 2898.4 22554.3
SubTrack++ (Ours) 627.6 1140.6 1593.2 1304.3 1517.6 16491.7
Appendix G Pre-Training LLama-Based Architectures
We pre-trained all six Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 10. Hyperparameters are selected based on reported values in GaLore paper [Zhao et al., 2024].
Table 10: Hyperparameters of pre-training Llama-based architectures.
60M 130M 350M 1B 3B 7B
Architectural Hidden 512 768 1024 2048 2560 4096
Parameters Intermediate 1376 2048 2736 5461 6848 11008
Heads 8 12 16 24 32 32
Layers 8 12 24 32 32 32
Shared Parameters Learning Rate 1e-3 1e-3 1e-3 1e-4 1e-4 1e-4
Batch Size 128 128 64 8 8 4
Gradient Accumulation 2 2 2 2 2 4
Iterations 10K
Gradient Clipping 1.0
Warmup Steps 1000
scale 0.25
dtype bfloat16
Low-Rank Optimizer Rank 128 256 256 512 512 1024
Methods Parameters Subspace Update Interval 200 200 200 200 200 500
SubTrack++ Step-Size 10
BAdam Parameters Block Swithch Interval 100
Switch Mode Random
Appendix H Subspace Tracking vs. SVD
Replacing SVD with Grassmannian-based subspace tracking, as opposed to the approach used in GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025], offers two key advantages: (1) Robustness: As shown in Figure 5, Grassmannian tracking better handles noise and abrupt shifts, preserving subspace quality under dynamic conditions. (2) Efficiency: It significantly reduces computational overhead (see Table 2 and Appendix D). Figure 6 illustrates the wall-time of the experiments in the ablation in Figure 3, highlighting substantial speedups.
As a result, SubTrack++ achieves SOTA-level performance with GaLore-like memory use, while being one of the fastest methods; clearly demonstrated in Figure LABEL:fig:modular and 6.
Figure 6: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ . Grassmannian tracking exceeds GaLore’s convergence, and significantly reduces wall-time, as also shown in Figure LABEL:fig:modular.